On average, each unit produced by the Kantor Company takes 0.90 worker hours and 0.02 hours of machine time. Furthermore, each worker and machine is available 160 hours a month. Use these planning values and the following sales order forecast to estimate

a) The number of worker hours and machine hours needed each month

b) The number of actual workers and machines needed each month (rounded to the nearest whole number)

Month Sales Forecast

October 44,000

November 52,000  

December 68,000

January 69,000

February 58,000

March 46,000
Answer #1

Given:

Worker hour per unit = 0.90 hours

Machine hour per unit = 0.02 hours

Total worker and machine hours available per month = 160 hours

(a)

One unit requires 0.90 worker hours. In the month of October, 44,000 units are forecasted. So the number of worker hours for 44,000 units will be 44,000*0.90. Similarly, for each month the calculation is done. For machine-hours, we will multiply by 0.02, as for each unit, 0.02 machine hours are required.

Month

Sales Forecast

Worker hours needed

Machine hours needed

October

44,000

44,000*0.90=39,600

44,000*0.02=880

November

52,000

52,000*0.90=46,800

52,000*0.02=1,040

December

68,000

68,000*0.90=61,200

68,000*0.02=1,360

January

69,000

69,000*0.90=62,100

69,000*0.02=1,380

February

58,000

58,000*0.90=52,200

58,000*0.02=1,160

March

46,000

46,000*0.90=41,400

46,000*0.02=920

(b)

The total worker and machine hours available are 160. So we need to divide available working hours by 160 to find out the number of workers and machines required per month.

Month

Sales Forecast

Number of Workers needed

Number of Machines needed

October

44,000

39,600/160=247.5=248

880/160=5.5=6

November

52,000

46,800/160=292.5=293

1,040/160=6.5=7

December

68,000

61,200/160=382.5=383

1,360/160=8.5=9

January

69,000

62,100/160=388.125=389

1,380/160=8.625=9

February

58,000

52,200/160=326.25=327

1,160/160=7.25=8

March

46,000

41,400/160=258.75=259

920/160=5.75=6